OpenStudy (anonymous):

Use complete the square, factoring or quadratic formula to complete the following. 2x^2-3=125

Parth (parthkohli):

\[\begin{aligned}2x^2 - 3 & = &125 \\ 2x^2 & = & 125 + 3\\ x^2& = & 128\div 2 \\ x^2 - 64& =&0 \\ (x + 8)(x - 8)&=&0\end{aligned}\]

OpenStudy (anonymous):

Which method did you use

Parth (parthkohli):

Look carefully through the steps.

OpenStudy (anonymous):

Completing the square?